OpenStudy (anonymous):

find the second derivative of e^(2x^3)

OpenStudy (anonymous):

\[12 e^{2 x^3} \left(x+3 x^4\right) \]

OpenStudy (anonymous):

lets say \[y=e^{2x^{3}}\] then first derivative: \[y^'=6x^2(e^{2x^{3}})\] and second derivative: (using product rule) \[y^''= 6x^2[6x^2(e^{2x^{3}})]+e^{2x^{3}}[12x]\] simplify, thats your second derivative

OpenStudy (anonymous):

\[\int\limits \left(\int\limits 12 e^{2 x^3} \left(x+3 x^4\right)dx\right) \, dx=e^{2 x^3} \]

OpenStudy (anonymous):

Thank you guys!!! I have one more question how do u find the second derivative of 3xe^x??

OpenStudy (anonymous):

\[3 e^x \text{Log}[e] (2+x \text{Log}[e]) \]

OpenStudy (anonymous):

\[\int\limits \left(\int\limits 3 e^x \text{Log}[e] (2+x \text{Log}[e])dx\right) \, dx=3 e^x x \]